A ball is released from a height of 20N. Calculate



a. The time it takes to fall



b. The velocity to which it hits the ground

1
Expert's answer
2021-11-16T09:53:10-0500

(a) We can find the time it takes to fall as follows:


"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\times20\ m}{9.8\ \dfrac{m}{s^2}}}=2.02\ s."

(b) We can find the velocity to which it hits the ground from the kinematic equation:


"v=v_0+gt=-9.8\ \dfrac{m}{s^2}\times2.02\ s=-19.8\ \dfrac{m}{s^2}."

The sign minus means that the velocity of the ball directed downward.